OpenStudy (anonymous):

What is the solution to the equation 14^ x + 3 = 21? x ≈ −2.133 x ≈ −1.846 x ≈ 3.867 x ≈ 4.154

OpenStudy (anonymous):

Oh, i have another o-o xD

OpenStudy (anonymous):

What transformation has changed the parent function f(x) = 3(2)x to its new appearance shown in the graph below?

OpenStudy (anonymous):

f(x) + 2 f(x) − 2 f(x + 2) f(x − 2)

OpenStudy (anonymous):

Another? Sorry, i just hate these kinds of graphs, they confuse meh.

OpenStudy (anonymous):

Yeah, i have a graphing calculator, but i might have too xD....What transformation has changed the parent function f(x) = (.5)x to its new appearance shown in the graph below?

OpenStudy (anonymous):

f(x) − 2 f(x + 2) f(x − 2) f(x) + 1

OpenStudy (anonymous):

Arent you just a ball of greatness. xD

OpenStudy (anonymous):

Which of the following is equivalent to log324? 0.4 0.424 0.125 2.5

OpenStudy (anonymous):

the 32 is small.

OpenStudy (anonymous):

Oh, Thanks ;o i must be annoying with all these questions xD.

OpenStudy (anonymous):

iven four functions, place them in order of their y-intercept, from highest to lowest.

OpenStudy (anonymous):

g(x), h(x), j(x), f(x) h(x), j(x), g(x), f(x) f(x), g(x), j(x), h(x) j(x), g(x), h(x), f(x)

OpenStudy (anonymous):

Oh, hold on o-o G(x) = j(x) 8(20)^x h(x)= Al is monitoring the decay of a population of fungi. It is reducing in half every four weeks. The population started at 5. j(x)= x-1,2,3 f(x)- 8-4-2

OpenStudy (anonymous):

c: Im just gonna cry. thanks for helping meh with all this. xD but yano, more to come. as you can see, math isnt my strongest subject.

OpenStudy (anonymous):

Given four functions, which one will have the smallest y-intercept? f(x)-Blake is tracking his savings account with an interest rate of 5% and a original deposit of $6 g(x)- x-0,1,2 f(x)-2,6,10 h(x)- See attachment. j(x)- j(x)=10(2)^x

OpenStudy (anonymous):

f(x) g(x) h(x) j(x)

OpenStudy (anonymous):

A heated piece of metal cools according to the function c(x) = (.5)x − 9, where x is measured in hours. A device is added that aids in cooling according to the function h(x) = −x − 2. What will be the temperature of the metal after five hours? −7° Celsius 6° Celsius 9° Celsius 16° Celsius

OpenStudy (anonymous):

Yes, yes it is.

OpenStudy (anonymous):

i got -7. But im second guessing.

OpenStudy (anonymous):

Given the function f(x) = log2(x + 6), find the value of f−1(3)?

OpenStudy (anonymous):

f−1(3) = 2 f−1(3) = 3 f−1(3) = 9 f−1(3) = 18

OpenStudy (anonymous):

i dont understand how your getting these, but okay xD

OpenStudy (anonymous):

What is the logarithmic function modeled by the following table? x-8,16,32 f(x)-3,4,5 f(x) = logx2 f(x) = log2x f(x) = 2 log10x f(x) = x log102

OpenStudy (anonymous):

A video game sets the points needed to reach the next level based on the function g(x) = 7(3)x − 1, where x is the current level. The hardest setting promises to multiply the points needed in each level according to the function h(x) = 4x. How many points will a player need on the hardest setting of level 4? 16 points 189 points 3024 points 12096 points

OpenStudy (anonymous):

x-1 is the exponent o-o

OpenStudy (anonymous):

The value of a dirt bike decreases by 20% each year. If you purchased this dirt bike today for $500, to the nearest dollar how much would the bike be worth 5 years later? $84 $119 $164 $222

OpenStudy (anonymous):

did you use A=Pe^rt?

OpenStudy (anonymous):

OHHH. Youre amazing ;3 i got a 42/46! WOOO. Im gonna be your fan c:

OpenStudy (anonymous):

Well thank you so much ! expecially for putting up with my Issues. i got the ones i did myself wrong, so no worries xD But THANK YOUU.